### v1

@@ Abstract
Quantum algorithms for matrix arithmetics, in particular solving linear system of equations (LSE), is a strongly studied subject in quantum computing. Most well-known might be the HHL algorithm [1], which promises a poly-logarithmic scaling in the system size, compared to the classical polynomial scaling [2]. However, this improvement is subject to a high sparsity and low condition number of the underlying system. Block encodings can be viewed as an generalization of this concept, with similar promises and requirements [3]. The application of these techniques has been studied in various fields, including training classical neural networks [4] and quantum reinforcement learning [5, 6]. However, executing these algorithms requires fault-tolerant and large-scale quantum devices.
@@ II Underlying Concept and Modifications
with all $A_{k}$ unitary and complex coefficients $c_{k}$.
@@ II-A Training with Modified Cost Function
In order to avoid unintentionally minimizing the norm of |$\ket{\psi}:=A\ket{V(\bm{\alpha})}$, it is common practice to employ a normalized version of the global cost function. Bravo-Prieto et al. suggest the formulation $C_{G}=\hat{C}_{G}/\expectationvalue{\psi|\psi}$ (see equation 5 in [8]). We instead propose
@@
which prevents division of the first summand of Equation 5 by $\expectationvalue{\psi|\psi}$ and therefore guarantees convergence of the cost function at a value of 00. With the same reasoning, we also implement an equivalent re-defined version of the normalized local cost function $C_{L}$ (see equation 6 and 7 in [8]).
@@ II-B Reduced Evaluation Overhead
where the second step employs $\beta_{kl}=\beta^{*}_{lk}$. As the terms $\beta_{kl}$ have to be evaluated individually – using e.g. the Hadamard test for the real part, and a modified version for the imaginary one – realizing Equation 7 instead of Equation 8 saves half the terms. Additionally, we exploit that the imaginary part of $\beta_{kl}$ can be ignored, if $c_{k}c_{l}^{*}\in\mathbb{R}$, leading to a further halving of the required Hadamard tests. An equivalent decomposition exploiting symmetries and real-valued coefficients is also implemented for the raw cost functions $\hat{C}_{G}$ and $\hat{C}_{L}$.
@@ III Framework Components and Usage
The proposed framework is based on PennyLane [13] and implements the routines proposed by Bravo-Prieto et al. [8], with the modifications discussed in Section II-A. The following section covers the details of the implementation, usage and potential extensions. An overview of the provided functionalities and respective constraints is provided in Table I. Additional usage information is also provided in the documentation of variational-lse-solver.
@@ III-A Modes of Loading System Matrix
One integral part of the routine is encoding the system matrix as unitary operations. While there is considerable research effort on the synthesis of circuits given an unitary, we approach this topic from a more abstracted view. Our framework allows to provide the decomposition of the system matrix in three different modes:
@@
Additionally, one can use a direct evaluation of the loss function described in Section III-D by only providing the potentially non-unitary matrix $A$.
@@ III-D Additional Evaluation Methods
We propose and implement additional techniques for the actual evaluation of the loss functions, an overview is provided in Table I. First of all, we explicitly realize the Hadamard-overlap test for evaluation of the local cost function, which was briefly mentioned in [8]. Second, we implement a version of the global cost function that exploits setups with coherent access to the unitary decomposition of $A$ [14, 13]. Last but not least, we provide a fully differentiable direct computation of the losses $C_{G}$ and $C_{L}$ following Equation 4. This method is not feasible on actual quantum hardware, but allows for fast validation and prototyping in simulation.
@@ IV Concluding Remarks
We introduce variational-lse-solver, a comprehensive library for research on variational linear system of equations (LSE) solvers. The framework allows researchers to abstract away technical details and focus on developing and refining end-to-end applications. We believe this tool will greatly benefit the community and contribute to the progress of quantum software development.
@@
The described library can be installed via pip install variational-lse-solver. The code with details on requirements and usage is also available at https://github.com/nicomeyer96/variational-lse-solver. Further information and data is available upon reasonable request.

### v2

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@
Quantum algorithms for matrix arithmetics, in particular solving linear system of equations (LSE), is a strongly studied subject in quantum computing. Most well-known might be the HHL algorithm [1], which promises a poly-logarithmic scaling in the system size, compared to the classical polynomial scaling [2]. However, this improvement is subject to a high sparsity and low condition number of the underlying system. Block encodings can be viewed as a generalization of this concept, with similar promises and requirements [3]. The application of these techniques has been studied in various fields, including training neural networks [4] and quantum reinforcement learning [5, 6]. However, executing these algorithms requires fault-tolerant and large-scale quantum devices.
@@ II Underlying Concept and Modifications
with all $A_{k}$ unitary and complex coefficients $c_{k}$. While decomposing a given matrix $A$ is typically a non-trivial task, there are applications in e.g. quantum chemistry [13] and quantum simulation [14], where one has (coherent) access to unitary realizations following Equation 2.
@@ II-A Training with Modified Cost Function
Intuitively, this can also be understood as maximizing the overlap between $A^{\dagger}\ket{b}$ and |$\ket{V(\bm{\alpha})}$, i.e. the ground-truth and the approximated solution of the linear system. In order to avoid unintentionally minimizing the norm of |$\ket{\psi}:=A\ket{V(\bm{\alpha})}$, it is common practice to employ a normalized version of the global cost function. Bravo-Prieto et al. suggest the formulation $C_{G}=\hat{C}_{G}/\expectationvalue{\psi|\psi}$ (see equation 5 in [8]). We instead propose
@@
which prevents division of the first summand of Equation 5 by $\expectationvalue{\psi|\psi}$ and therefore guarantees convergence of the cost function at a value of 00. With the same reasoning, we also implement an equivalent re-defined version of the normalized local cost function $C_{L}$ (see equations 6 and 7 in [8]).
@@ II-B Reduced Evaluation Overhead
where the second step employs $\beta_{kl}=\beta^{*}_{lk}$. As the terms $\beta_{kl}$ have to be evaluated individually – using e.g. the Hadamard test for the real part, and a modified version for the imaginary one – realizing Equation 7 instead of Equation 8 saves half the terms. Additionally, we exploit that the imaginary part of $\beta_{kl}$ can be ignored if $c_{k}c_{l}^{*}\in\mathbb{R}$, leading to a further halving of the required Hadamard tests. An equivalent decomposition exploiting symmetries and real-valued coefficients is also implemented for the raw cost functions $\hat{C}_{G}$ and $\hat{C}_{L}$.
@@ III Framework Components and Usage
The proposed framework is based on PennyLane [15] and implements the routines proposed by Bravo-Prieto et al. [8], with the modifications discussed in Section II-A. For training the parameters of the underlying variational circuit, by default an Adam optimizer [16] with a learning rate of $0.01$ is employed. We found this to be the most stable choice for our experiments, but also ensure customizability for varying setups. Furthermore, the library allows for defining early-stopping criteria, based on the operational meaning of the cost function [8].
The following sections cover the details of the implementation, usage, and potential extensions. An overview of the provided functionalities and respective constraints is provided in Table I. Additional usage information is also provided in the documentation of variational-lse-solver.
@@ III-A Modes of Loading System Matrix
One integral part of the routine is encoding the system matrix as unitary operations. While there is research effort on the synthesis of circuits given an unitary [17], we approach this topic from a more abstracted view. Our framework allows to provide the decomposition of the system matrix in three different modes:
@@
Additionally, one can employ direct evaluation of the loss function described in Section III-D by only providing the potentially non-unitary matrix $A$. However, this only is feasible in simulation, as it sidesteps the unitary decomposition following Equation 2.
@@ III-B Usage Example for Recreating Simple Experiment
An example of work integrating the presented library in a more complex routine can be found in Meyer et al. [18]. More concretely, there the variational algorithm is used to solve linear systems underlying a reinforcement learning setup.
@@ III-D Additional Evaluation Methods
We propose and implement additional techniques for the actual evaluation of the loss functions, an overview is provided in Table I. First of all, we explicitly realize the Hadamard-overlap test for evaluation of the local cost function, which was briefly mentioned in [8]. Second, we implement a version of the global cost function that exploits setups with coherent access to the unitary realization of $A$ [19, 15], thus avoiding the need for explicit decomposition and circuit synthesis. Last but not least, we provide a fully differentiable direct computation of the losses $C_{G}$ and $C_{L}$ following Equation 4. This method is not feasible on actual quantum hardware but allows for fast validation and prototyping in simulation.
@@ IV Concluding Remarks
We introduce variational-lse-solver, a comprehensive library for research on variational linear system of equations (LSE) solvers. It has to be noted, that it is currently still under debate, in which scenarios, and under which conditions, such techniques are applicable and beneficial. Especially the decomposition of the system matrix, and subsequent realization of the unitary matrices as quantum circuits, requires further attention. To allow for progress toward identifying potentials and further bottlenecks, our framework allows researchers to abstract away those technical details. Due to this, believe this tool will greatly benefit the community and contribute to the progress of quantum software development.
@@
The presented python library can be easily installed via pip install variational-lse-solver. The implementation, with further details on requirements and usage instructions, is also available at https://github.com/nicomeyer96/variational-lse-solver Further information and data is available upon reasonable request.